Cash and Equivalents = $14,000: Accounts Receivable= $28,000: Accounts Payable = $18,000 Short term debt= $4000: Inventories= $16,000, Other current liabilities= $12,000: and other current assets = $8,000. What is networking capital? A) -$16,800 B) -$32,000 C) $32,000 D) $42,000
Answer #1

Total current assets=Cash and equivalents+AR+Inventory+Other current assets

(14000+28000+16000+8000)=$66000

Total current liabilities=AP+Short term debt+Other current liabilities

=(18000+4000+12000)=$34000

Net working capital=Total current assets-Total current liabilities

=(66000-34000)

=$32,000.